OpenStudy (anonymous):

NOTE :THIS IS NOT A QUESTION,THIS IS A TUTORIAL. PH value & PoH value ==============

OpenStudy (anonymous):

PH Value : __________ *First of all ,what Ph Value mean ?? ===> pH is a measure of the hydrogen ion concentration, [H+] . *what is the formula which used to calculate the pH value ?? ===> pH is calculated using the following formula: pH = -log10[H+] . *_* Here is some examples about PH value and how to find it : ============================================================== Example 1: ___________ Find the pH of a 0.2mol L^-1 (0.2M) solution of HCl . Ans: 1-Write the balanced equation for the dissociation of the acid HCl -----> H+(aq) + Cl-(aq) 2- Use the equation to find the [H+]: 0.2 mol L- HCl produces 0.2 mol L-1 H+ since HCl is a strong acid that fully dissociates 3- Calculate pH: pH = -log10[H+] pH = -log10[0.2] = 0.7 Example 2: __________ Find the pH of a 0.2 mol L^-1 (0.2M) solution of H2SO4 . Ans: 1-Write the balanced equation for the dissociation of the acid H2SO4 -----> 2H+(aq) + SO42-(aq) 2- Use the equation to find the [H+]: 0.2 mol L-1 H2SO4 produces 2 x 0.2 = 0.4 mol L^-1 H+ since H2SO4 is a strong acid that fully dissociates 3- Calculate pH: pH = -log10[H+] pH = -log10[0.4] = 0.4 ****-A Tip: Hydrogen ion concentration, [H+], can be calculated using the following formula: [H+] = 10-pH . Example 3: ____________ *Find the [H+] of a nitric acid solution with a pH of 3.0 . Ans: pH= 3.0 [H+] = 10-pH [H+] = 10-3.0 = 0.001mol L^-1 You can check this answer by using the calculated value [H+] in the equation for pH to make sure you arrive at the original pH pH = -log10[H+] pH = -log10[0.001] = 3 We get the same value for pH using the calculated value for [H+], so the calculated value for [H+] is correct. ______________________________________________________________________________________________________________________________ Now Lets Move To POH Value: =========================== *-First of all ,What is POH value ? ====>pOH is a measure of the hydroxide ion concentration, [OH-]. *-what is the formula which used to calculate the poH value ?? ===>pOH is calculated using the following formula: pOH = -log10[OH-]. Now here is some Examples: =_=_=_=_==_=_=_=_==_=_=_=_= Example 1: _________ Find the pOH of a 0.1mol L^-1 (0.1M) solution of NaOH . Ans: 1-Write the balanced equation for the dissociation of the alkali NaOH -----> OH-(aq) + Na+(aq) 2- Use the equation to find the [OH-]: 0.1 mol L-1 NaOH produces 0.1 mol L^-1 OH- since NaOH is a strong alkali that fully dissociates 3- Calculate pOH: pOH = -log10[OH-] pOH = -log10[0.1] = 1 Example 2: __________ Find the pOH of a 0.1mol L^-1 (0.1M) solution of Ba(OH)2 . Ans: 1-Write the balanced equation for the dissociation of the alkali: Ba(OH)2 -----> 2OH-(aq) + Ba2+(aq) 2- Use the equation to find the [OH-]: 0.1mol L^-1 Ba(OH)2 produces 2 x 0.1 = 0.2 mol L^-1 OH- since Ba(OH)2 is a strong alkali that fully dissociates 3- Calculate pOH: pOH = -log10[OH-] pOH = -log10[0.2] = 0.7 *_* A TIP : Hydroxide ion concentration, [OH-], can be calculated using the following formula: [OH-] = 10-pOH . Example 3: __________ Find the [OH-] of a sodium hydroxide solution with a pOH of 1 pOH = 1 [OH-] = 10-pOH [OH-] = 10-1 = 0.1 mol L^-1 Ans: ==== pOH = 1 [OH-] = 10-pOH [OH-] = 10-1 = 0.1 mol L^-1 You can check this answer by using the calculated value [OH-] in the equation for pOH to make sure you arrive at the original pOH pOH = -log10[OH-]. ===================================================================================== If U have any questions ,please let me Know :)

OpenStudy (anonymous):

nicely done really! now just save yourself link to this "question" as i did so you can share leater upon question :D (and just from begining pH is measurement of activity (a) of H+ (H3O+) ions because pH doesnt have unit and you cant log unit... but great work nevertheless)

OpenStudy (anonymous):

Ty alot @Kryten . :)

OpenStudy (unklerhaukus):

this tutorial is very good. can i ask what the ranges (in practice and theoretical) for pH/pOH are

OpenStudy (zepp):

Nice!

OpenStudy (anonymous):

@zepp :Ty :)

OpenStudy (anonymous):

@UnkleRhaukus ranges for pH and pOH practicaly can go a lot further than 0-14, you have super acids which have pH of about -30 and bases with pH about 20. my sudgestion for you is to type super acid in google cause momentarily i dont have time to explain it... :) (hope you are not mad)

OpenStudy (anonymous):

dayum. this kid is smart ;P